Find the upward flux of F vector=<x,y,z> across: a. x^2+y^2+z^2=1, z\geq0 and b. z=1-x^2-y^2, z\geq0.

Please be detail thanks.
Answer #1

F <x, y, z) (a) x² + y + z =1, 27,0 1-x-y? z = Z- 51-2²-y- fray) = -xt-yl fx- da (-2x) 5-a²-4 251-2-y? a six-ye fy - y 251 -